# Spherical solutions to the Klein-Gordon equation in the expanding universe

### v1

@@ Abstract
In this study, we produce an explicit wave function of the particles (fields) emitted to the FLRW universe with the scale factor generated by the de Sitter universe. We consider models of scalar fields with spherical symmetry at the initial time. They can be generated, for instance, by a pionic atom, exotic atom, or gravitational field (see, e.g., [8, 14, 5, 26]). In the expanding universe, the particles obey the equation for the free particles. A purpose of this article is a derivation of the explicit form of exact solutions of the spherical waves in the curved spacetime. The models illustrate an instantaneous transition from the initial spacetime to the non-static FLRW universe with a de Sitter scale factor.
@@
$\cases{\displaystyle\Box_{g}\Phi+\frac{m_{n}^{2}c^{2}}{\hbar^{2}}\Phi=0\,,\cr\displaystyle\Phi(x,0)=F(r,0)Y_{\ell m}(\theta,\phi)\,,\quad\Phi_{t}(x,0)=F_{t}(r,0)Y_{\ell m}(\theta,\phi)\,,}$
@@
$\square_{g}\Phi=\frac{1}{\sqrt{|g|}}\frac{\partial}{\partial x^{i}}\left(\sqrt{|g|}g^{ik}\frac{\partial\Phi}{\partial x^{k}}\right).$
We are interested in the explicit form of the spherical field in the de Sitter FLRW spacetime, which is a solution of the problem
@@
Here, $Y_{lm}(\theta,{\phi})$ is a spherical harmonic, $\ell=0,1,\ldots$, $m=0,\pm 1,\pm 2\ldots$, $|m|\leq\ell$. In particular, the initial state can be given by a wave function $\mbox{\rm Exp}\left(-i\frac{E_{n}t}{\hbar}\right)\frac{R_{n\ell}(r)}{r}Y_{\ell m}(\theta,\phi)$ of the pionic atom (see, e.g., [8]). We will use a system of units in which the speed of light $c=1$.
@@
We often omit the parameter $M$ in $K_{0}(z,t;M)$ and $K_{1}(z,t;M)$ when it does not lead to confusion. Denote $\phi(t):=(1-e^{-Ht})/H$. The main results of this study, which are obtained by the integral transform approach (ITA) [21, 24], are the following two theorems.
@@
It should be noted here that for negative integers $a$, the hypergeometric function $F(a,b;c;z)$ is a polynomial of degree a-a (see, e.g., [3]).
In the next theorem, by imposing conditions on the behavior of the functions $F_{0}=F_{0}(r)$ and $F_{1}=F_{1}(r)$ for large $r$ one can appeal to the Hankel transform and write the representation of the field in the form that is more amendable for the $L^{2}({\mathbb{R}}^{3})$-estimates. In particular, this applies to the wave function of the pionic atom that instantly arrives in the de Sitter FLRW spacetime, because that wave function has an exponential decay at infinity of spatial coordinates (see, e.g., [8]).
@@
Below, we provide some examples of such an initial state $(\Phi(x,0),\Phi_{t}(x,0))$ of the field. In particular, it can be a wave function of a pionic atom in the Minkowski spacetime (see, e.g., [8, 10]), or a scalar field in the matter- or radiation-dominated or de Sitter universe with a different Hubble constant $H_{0}\not=H$ and different mass. In exotic atoms, one of the electrons has been replaced by a negatively charged particle such as a muon, K-meson, pion, $\Sigma$-hyperon, etc. For further details on the physical interest in such systems, see, for example, [8, 9, 10, 13]. A pionic atom is a hydrogen-like system in which an electron replaced by a spinless relativistic particle pion. A pion is an example of a particle with electromagnetic interactions that obeys the Klein-Gordon equation. In fact, by pionic atom, one can test the validity of the Klein-Gordon equation (see, e.g., [10, Sec.18.4]). In (1.2), the pion has instantly left the atom, the Coulomb field is no longer in action, and spacetime has instantly become inflationary. Thus, the free pion propagates in the expanding universe and carries information about its previous state. Another possible application can be to the problem of the existence of quasinormal modes on de Sitter space (see [12]).
Although spherically symmetric solutions for the Klein-Gordon and Dirac equations in the Minkowski spacetime have been known for a long time (see, e.g., [4, §24], [7], [15], and references therein), explicit exact formulas in the curved spacetimes are known only for a few special cases (see, e.g., [2, 6, 18, 25]). We believe that the obtained explicit formulas of Theorem 1.2 can be useful, in particular, in the study of caustics for spherical waves in curved spacetime (see, e.g., [7]). In the forthcoming paper, the representation formulas of Theorems 1.1, 1.2 will be used for the derivation of the exact spherical solutions to the spin-$\frac{1}{2}$ fields in the expanding universe.
@@ 2.1 The general solution approach
We look for the solution $F=F(r,t)$ of equation (2.3) satisfying (2.4). For example, if the initial state $F_{0}(r)$ is the the wave function of the pionic atom [8, pp.188-189], [10, Sec.18.4], and [14, §26, Sec 2] in the Minkowski space, then
@@
If $\ell=0$, while $Z=1$ and $Z=2$, then $\mu-\frac{1}{2}=-0.0000532822$ and $\mu-\frac{1}{2}=-0.000213163$, respectively.
@@
where $a:=\mu-\frac{1}{2}+n-\ell-1+k$. The last integral can be evaluated in terms of the incomplete gamma function ([3, Ch.9]):
@@ 2.3 Solution by Hankel transform
with $\nu>-1/2$ ([16, Sec.11.5.1], [19, Sec.14.3], [27, Sec.14.3]). Let $f(r)$ be a function defined for $r>0$. The inversion formula yields
@@
$\Phi(r,\theta,\phi,t)$ $=$ $Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda t\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda$
$-Y_{\ell m}(\theta,\phi)\frac{1}{\sqrt{r}}\int_{0}^{t}\Bigg{\{}\frac{m_{0}t}{\sqrt{t^{2}-\tau^{2}}}J_{1}\left(m_{0}\sqrt{t^{2}-\tau^{2}}\right)-J_{0}\left(m_{0}\sqrt{t^{2}-\tau^{2}}\right)\Bigg{\}}$
$\times\Bigg{\{}\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{1}(\rho)J_{\ell+\frac{1}{2}}(\rho\lambda)\rho^{3/2}\,d\rho\right)\cos\left(\lambda\tau\right)J_{\ell+\frac{1}{2}}(r\lambda)\lambda\,d\lambda\Bigg{\}}\,d\tau$
@@
If $m_{n}=\sqrt{2}H\hbar$, then
$\left|\psi(x,t)-e^{-Ht}v_{\psi_{0}}(x,\phi(t))\right|$ $\leq$ $C(r)e^{-Ht}\left|Y_{\ell m}(\theta,\phi)\right|\,.$

### v2

@@ Abstract
In this study, we produce an explicit wave function of the particles (fields) emitted to the Friedmann-Lemaître-Robertson-Walker (FLRW) universe with the scale factor generated by the de Sitter space. We consider models of scalar fields with spherical symmetry at the initial time. They can be generated, for instance, by a pionic atom, exotic atom, or gravitational field (see, e.g., [1, 2, 3, 4]). In the expanding universe, the particles obey the equation for the free particles. The purpose of this article is to derive the explicit form of exact solutions of the spherical waves in curved spacetime. The models illustrate a transition from the initial spacetime to the non-static FLRW universe with a de Sitter scale factor.
@@
$\cases{\displaystyle\Box_{g}\Phi-\frac{m_{n}^{2}c^{2}}{\hbar^{2}}\Phi=0\,,\cr\displaystyle\Phi(x,0)=F(r,0)Y_{\ell m}(\theta,\phi)\,,\quad\Phi_{t}(x,0)=F_{t}(r,0)Y_{\ell m}(\theta,\phi)\,,}$
@@
$\square_{g}\Phi=\frac{1}{\sqrt{|g|}}\frac{\partial}{\partial x^{\alpha}}\left(\sqrt{|g|}g^{\alpha\beta}\frac{\partial\Phi}{\partial x^{\beta}}\right),$
where $\alpha$ and $\beta$ are running from 0 to $3$. We are interested in the explicit form of the spherical field in the de Sitter FLRW spacetime, which is a solution of the problem
@@
Here, $Y_{lm}(\theta,{\phi})$ is a spherical harmonic, $\ell=0,1,\ldots$, $m=0,\pm 1,\pm 2\ldots$, $|m|\leq\ell$. In particular, the initial data $\Phi(x,0)$ and $\Phi_{t}(x,0)$ can be given by a wave function of the pionic atom in Minkowski space, that is, by the functions $\frac{R_{n\ell}(r)}{r}Y_{\ell m}(\theta,\phi)$ and $\left(-i\frac{E_{n}}{\hbar}\right)\frac{R_{n\ell}(r)}{r}Y_{\ell m}(\theta,\phi)$, respectively (see, e.g., [2]). We will use a system of units in which the speed of light $c=1$.
@@
We often omit the parameter $M$ in $K_{0}(z,t;M)$ and $K_{1}(z,t;M)$ when it does not lead to confusion. Denote $\phi(t):=(1-e^{-Ht})/H$ a conformal time. Note: the scalar curvature $R$ is encoded in these kernels via the Hubble constant $H=\sqrt{R/12}$. The main results of this study, which are obtained by the Integral Transform Approach (ITA) [6, 7], are the following two theorems.
@@
In the case of $m_{n}^{2}=2H^{2}\hbar^{2}$, the Klein-Gordon equation becomes the conformally invariant wave equation. It should be noted here that for negative integers $a$, the hypergeometric function $F(a,b;c;z)$ is a polynomial of degree a-a (see, e.g., [5]).
In the next theorem, by imposing conditions on the behavior of the functions $F_{0}=F_{0}(r)$ and $F_{1}=F_{1}(r)$ for large $r$ one can appeal to the Hankel transform and write the representation of the field in the form that is more amenable for the $L^{2}({\mathbb{R}}^{3})$-estimates. In particular, this applies to the wave function of the pionic atom that arrives in the de Sitter FLRW spacetime, because that wave function has an exponential decay at infinity of spatial coordinates (see, e.g., [2]).
@@
In Corollary 3.2, we prove the exponential decay of the solution’s tail in the case where the radial part of the initial function is generated by the wave function of the pionic atom. Moreover, it is also shown that the obtained decay is optimal.
Below, we provide some examples of such an initial state $(\Phi(x,0),\Phi_{t}(x,0))$ of the field. In particular, it can be a wave function of a pionic atom in the Minkowski spacetime (see, e.g., [2, 8]), or a scalar field in the matter- or radiation-dominated or de Sitter universe with a different Hubble constant $H_{0}\not=H$ and different mass. In exotic atoms, one of the electrons has been replaced by a negatively charged particle such as a muon, K-meson, pion, $\Sigma$-hyperon, etc. For further details on the physical interest in such systems, see, for example, [2, 8, 9, 10]. A pionic atom is a hydrogen-like system in which an electron replaced by a spinless relativistic particle pion. A pion is an example of a particle with electromagnetic interactions that obeys the Klein-Gordon equation. In fact, by pionic atom, one can test the validity of the Klein-Gordon equation (see, e.g., [8, Sec.18.4]). In (1.2), the pion has left the atom, the Coulomb field is no longer in action, and spacetime has become inflationary. Thus, the free pion propagates in the expanding universe and carries information about its previous state.
Although spherically symmetric solutions for the Klein-Gordon and Dirac equations in the Minkowski spacetime have been known for a long time (see, e.g., [11, §24], [12], [13], and references therein), explicit exact formulas in the curved spacetimes are known only for a few special cases (see, e.g., [14, 15, 16, 17]). We believe that the obtained explicit formulas of Theorem 1.2 can be useful, in particular, in the study of caustics for spherical waves in curved spacetime (see, e.g., [12]). In the forthcoming paper, the representation formulas of Theorems 1.1, 1.2 will be used for the derivation of the exact spherical solutions to the spin-$\frac{1}{2}$ fields in the expanding universe.
Another possible application of our formulas can be to the problem of the existence of quasinormal modes (QNMs) on a non-static FLRW universe with a de Sitter scale factor. The existence of QNMs on de Sitter spacetime has been much studied in the literature; a detailed analysis of the literature on the spherical waves and QNMs has been given in [18]. In [18], it has been demonstrated that QNMs exist for massive scalar fields on static de Sitter space in all spacetime dimensions greater than one and for all complex-valued scalar field masses. The Taylor expansion with the recursion relation for the coefficients has been presented for calculating of QNMs and their corresponding mode solutions. The proof of the convergence of the series has been done for the analytic data.
The study of the spherically symmetric solutions of the wave equation in the de Sitter space in the Bondi coordinates was carried out in [1]. In those coordinates, the problem was reduced to an integro-differential equation with the initial data on a light cone. The main results state existence and uniqueness, as well as decay of the solution in the Bondi coordinates and its limit at infinity.
@@ 2.1 The general solution approach
We look for the solution $F=F(r,t)$ of equation (2.3) satisfying (2.4). For example, if $F_{0}(r)$ is generated by the wave function of the pionic atom [2, pp.188-189], [8, Sec.18.4], and [3, §26, Sec 2] in the Minkowski space, then
@@
If $\ell=0$, while the atomic number $Z=1$ and $Z=2$, then $\mu-\frac{1}{2}=-0.0000532822$ and $\mu-\frac{1}{2}=-0.000213163$, respectively.
@@
Here $a_{k,\alpha,\beta}^{(\ell)}$ and $b_{k,\alpha,\beta}^{(\ell)}$ are some suitable coefficients.
@@
where $a:=\mu-\frac{1}{2}+n-\ell-1+k$. The last integral can be evaluated in terms of the incomplete gamma function ([5, Ch.9]):
@@ 2.3 Solution by Hankel transform
with $\nu>-1/2$ ([24, Sec.11.5.1], [25, Sec.14.3], [26, Sec.14.3]). Let $f(r)$ be a function defined for $r>0$. The inversion formula yields
@@
$\Phi(r,\theta,\phi,t)$ $=$ $Y_{\ell m}(\theta,\phi)\frac{1}{r}\Bigg\{\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{2}d\rho\right)\cos(\lambda t)J_{\ell+\frac{1}{2}}(r\lambda)\,\lambda d\lambda$
$-\int_{0}^{t}\frac{mt}{\sqrt{t^{2}-\tau^{2}}}J_{1}\left(m\sqrt{t^{2}-\tau^{2}}\right)\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{0}(\rho)J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{2}d\rho\right)\cos(\lambda\tau)J_{\ell+\frac{1}{2}}(r\lambda)\,\lambda d\lambda\,d\tau$
$+\int_{0}^{t}J_{0}\left(m\sqrt{t^{2}-\tau^{2}}\right)\int_{0}^{\infty}\left(\int_{0}^{\infty}F_{1}(\rho)J_{\ell+\frac{1}{2}}(\lambda\rho)\,\rho^{2}d\rho\right)\cos(\lambda\tau)J_{\ell+\frac{1}{2}}(r\lambda)\,\lambda d\lambda\,d\tau\Bigg\}$
@@
Thus, one can regard ITA as an analytical mechanism that, from the massless virtual field in the Minkowski space, generates a massive field in the curved spacetime. The integral transform condenses the evolution of that virtual scalar wave in the conformal time from its entry into the expanding universe to the moment of observation.
@@
$\leq$ $C(r)e^{-\frac{3}{2}Ht}\left|Y_{\ell m}(\theta,\phi)\right|\cases{e^{\frac{1}{2}Ht}\quad\mbox{\rm if}\quad 3H\hbar/2>m_{n}\geq\sqrt{2}H\hbar\,,\cr e^{t\sqrt{\frac{9}{4}H^{2}-\frac{m_{n}^{2}}{\hbar^{2}}}}\quad\mbox{\rm if}\quad m_{n}<\sqrt{2}H\hbar\,.}$
@@
Acknowledgments
The author is indebted to the reviewers for the valuable remarks and suggestions, which improved the text.